Two tablespoons of medication is equal to a. 15 cc b. 1 mL c. 20...

  1.        Two tablespoons of medication is equal to

a.

15 cc

b.

1 mL

c.

20 c

d.

30 mL

30 ans

  1.       If a patient is prescribed a medication that is available in 25-mg tablets and is directed to take 75 mg daily, how many tablets will be given for a 30-day supply?

a.

55

b.

90

c.

100

d.

110

  1.         An infant is 60 cm long. Convert this to inches.

a.

24 in

b.

29 in

c.

34 in

d.

50 in

  1.           The time 2017h can be identified as

a.

8:17 PM

b.

8:17 AM

c.

10:17 PM

d.

10:17 AM

  1.        The time 230am can be identified as
  1. 230h
  2. 0230
  3. 1430h
  4. 0230h
  1.         Zithromax® oral suspension is available in dropper form. When it is properly mixed, there are 200 mg per 5 mL. How many milligrams are taken if the patient took 2 teaspoons of the properly mixed suspension?

a.

100

b.

200

c.

300

d.

400

  1.         A medication with two active ingredients is ordered as 30 mg/1200 mg per tablet. The available tablets are 15 mg/600 mg. How many tablets would be given?

a.

1/2

b.

1

c.

1 1/2

d.

2

  1.        A medication is ordered as 6 mg tablets. Available is a 4 mg tablet. The number of tablets to be given is

a.

1/2

b.

1

c.

1 1/2

d.

2

  1.        A doctor prescribes 10 mg of cough syrup. Available is 5 mg/5 mL. How many teaspoons would you give?

a.

1/2

b.

1

c.

1 1/2

d.

2

  1.        The time 2045h can be identified as
  1. 8:45pm
  2. 2:45am
  3. 9:45pm
  4. 6:45pm
Answer #1

Answer :(d) two table spoon= 30 ML

15 cc= 15 ml

1 cc=1 ml

  • Available dose= 25 mg tablets

Ordered dose= 75 mg = 3 tablets ( 75/25 = 3 )

ie,3 tablets per day.

Number of tablets for 30 days= 30* 3= 90

Answer (b) 90 tablets for 30 days.

60 cm= x inches

1 cm= 0.39 inches= 0.4 inches

So, 60 cm= 60*0.4= 24 inches

. answer (a)

2017 hour= 20 hours 17 minutes

Since 01 pm = 13 hours

02 pm= 14 hours

03 pm= 15 hours

04 pm= 16 hours

5 pm= 17 hours

Likewise, 20 hours= 08 PM

So, 201 hours= 8.17 PM

Answer (a)

230 AM =02 hours 30 minutes morning

1430 h=. 2.30pm

ie, 0230 h

Answer (d)

1 teaspoon= 4.9 ml

2 teaspoon=. 9.8 ml(2*4.9)

5 ml contains 200 mg

So, 1 ml= 200 mg/ 5 ml= 40 mg

Therefore, 9.8 ml= x mg

= 9.8 ml* 40 mg= 392 mg approximately 400 mg

Answer ( d)

Ordered dose= 30 mg/1200 mg

Available dose= 15 mg/ 600 mg per tablet.

Ordered dose is twice of available dose

So ,2 tablets can provide ordered dose.

Answer (d)

Ordered dose= 6 mg tablet

Available dose= 4 mg tablet

No of tablet to be administered= 6 mg/ 4 mg = 1.5 tablet

Answer ( c)

1 teaspoon= 4.929 ml approximately 5 ml

So ,10 ml= 2 teaspoon

Ordered dose= 10 mg

5 ml = 5 mg

So 1 ml= 1 mg

So 10 ml= 10 mg

Answer (d) 2 teaspoon

2045 h

1300h= 1 pm

1400h= 2 pm

Likewise, 2000h = 8pm

So, 2045= 08hours 45 minutes

= 8.45 PM.

Answer (a)